The curve 𝑦=βˆ’π‘₯3+3π‘₯2+6π‘₯βˆ’8 cuts the π‘₯-axis at π‘₯=βˆ’2,π‘₯=1 and π‘₯=4.

a. Sketch the curve, showing clearly the intersection with the coordinate axes.

b. Differentiate 𝑦=βˆ’π‘₯3+3π‘₯2+6π‘₯βˆ’8

c. Show that the tangents to the curve at π‘₯=βˆ’2 and π‘₯=4 are parallel.


1
Expert's answer
2022-05-31T12:50:34-0400

a.



b.


"y'=(-x^3+3x^2+6x-8)'=-3x^2+6x+6"

c.


"slope_1=m_1=y'(-2)"

"=-3(-2)^2+6(-2)+6=-18"


"slope_2=m_2=y'(4)"

"=-3(4)^2+6(4)+6=-18"

Since "m_1=-18=m_2," then the tangents to the curve at "\ud835\udc65=\u22122" and "\ud835\udc65=4" are parallel.